OpenStudy (fanduekisses):

Particle Motion: A particle moves along the x-axis so that its position at any time t is given by x(t)= t^4-6t^2-3.

OpenStudy (fanduekisses):

Which of the following best describes the motion of the particle for 0<t<1? A) Moving to the right and speeding up B) Moving to the left and speeding up C)Moving to the right and slowing down. D) Moving to the left at a constant speed E)Moving to the left and slowing down

OpenStudy (anonymous):

first derivative second derivative check the signs of each on that interval

OpenStudy (fanduekisses):

So I'll be looking for the velocity and acceleration?

OpenStudy (fanduekisses):

ok :)

OpenStudy (anonymous):

yeah exactly if velocity is positive, forward if the acceleration is positive, speeding up etc

OpenStudy (anonymous):

guess i meant "moving right"

OpenStudy (fanduekisses):

ok so v(t)=4t^3-12t a(t)=12t^2-12

OpenStudy (fanduekisses):

Wait, don't I have to factor ?

OpenStudy (fanduekisses):

find what t equals and then plug in?

OpenStudy (campbell_st):

you need to factor... \[x'(t) = 4t(t^2 - 3)\] set x'(t) = 0 and find the stationary points... \[x''(t) = 12(t^2 - 1)\]

OpenStudy (fanduekisses):

t= \[\pm \sqrt{1}\]

OpenStudy (fanduekisses):

So the particle is at rest when t=+-sqrt(1)

OpenStudy (campbell_st):

so the velocity is zero at \[t = 0 ~~~and~~~ t = \sqrt{3}\] so substitute 0.5 into the velocity and you get a negative value for velocity I did this because it is on the given interval... so the slope of the tangent is negative between 0 and 1 so the particle is moving left... now you can look at the acceleration

OpenStudy (campbell_st):

well you can't have negative time... so the acceleration is zero at t = 1 and the acceleration is -12 at t = 0 so acceleration is going from -12 at t = 0 to 0 at t = 1 what do you think happened...?

OpenStudy (fanduekisses):

Hmmm so the particle stayed at constant velocity when t=1?

OpenStudy (campbell_st):

well I think the particle is moving left and slowing.... that's my best guess.

OpenStudy (fanduekisses):

so t= time right? how come the acceleration is -12 at zero seconds (time)?

OpenStudy (campbell_st):

well the velocity is zero, and by the definitions of velocity and acceleration, this is where the acceleration is a maximum... remember t = 0 is where velocity is 0.

OpenStudy (fanduekisses):

oooooh yeah lol

OpenStudy (fanduekisses):

Well, thanks so much for the help :D I gotta go now :)

OpenStudy (campbell_st):

here is a summary that may explain things https://apmathteacher.wikispaces.com/file/view/Particle+Motion.pdf